# FastSet: Parallel Claim Settlement

### v1

@@ 3 Formalization and Correctness
Unfortunately, the concurrency model underlying FastSet differs from classic (replica-based) distributed systems models like those in [21]. In distributed systems models, a number of communicating nodes perform a computation together. Accesses to shared objects impose certain constraints on the distributed computation, such as all the reads seeing the latest write value, but are also opportunities for nodes to share causality information. In FastSet, validators do not initiate any computation nor exchange any information with each other, neither directly nor indirectly through clients, and there is no causal dependence shared through updates (nor defined or calculated). Clients drive the protocol by broadcasting claims that may update the validator’s states. The only means of communication in FastSet, if it can even be called so, is that a client claim reaching a validator may be invalid for that validator, in which case the validator either discards it (in Step 3) or postpones it (in Step 6). An abstraction similar to our weak independence was introduced in [21] (Definition 2.6, Commutativity) and used to prove the SEC of their model, but we were not able to directly use it or their SEC theorem in our formalization.
@@
Validators will never get stuck due to unfortante interleaving orderings of claims. That is, if progress has been made by any other validators for any clients, then the validator in question can also make progress.
@@ 3.2 Weak Independence
Weak equivalence is reminiscent of the classic notion of “weak equality” in partial algebra (see, e.g., [4] for a survey): two claim sequences are weakly equivalent iff they result in the same state whenever they are both valid. What may be counter-intuitive at first sight is that an invalid claim sequence is weakly equivalent to any other claim sequence, whether valid or not – this meaning of “weak” is, however, well-established in the literature. This implies, in particular, that weak equivalence is not transitive: if $\gamma\Downarrow\gamma^{\prime}$ and $\gamma^{\prime}\Downarrow\gamma^{\prime\prime}$ then there is nothing to prevent the existence of a state $s$ such that ${\gamma}\!\downarrow_{s}$ and ${\gamma^{\prime\prime}}\!\downarrow_{s}$, but ${\gamma^{\prime}}\!\ndownarrow_{s}$ and $\llbracket{\gamma}\rrbracket{s}\neq\llbracket{\gamma^{\prime\prime}}\rrbracket{s}$. For example, $s$ can be a state where Alice has 11 tokens, $\gamma$ and $\gamma^{\prime\prime}$ payment claims by which Alice pays Bob 10 tokens and 11 tokens, respectively, and $\gamma^{\prime}$ some invalid claim (e.g., Alice pays Charlie 100 tokens).
@@
The reverse is not true. For example, if the two claims are same mutex acquires then each of them is valid independently, but they are not valid together regardless of the order. So $c\,c^{\prime}\Downarrow c^{\prime}\,c$ vacuously holds, but $c\parallel c^{\prime}$ does not hold.
@@ 3.3 Interleavings
Interleaving equivalence is reminiscent of actor/thread interleaving in concurrent programming, each address acting as an actor/thread executing a sequence of commands/claims. An interleaving cut $\gamma\sqsubseteq\gamma^{\prime}$ captures the intuition that $\gamma$ and $\gamma^{\prime}$ are two attempts to execute determined sequences of claims for each address, where $\gamma$ was cut short or didn’t advance as much as $\gamma^{\prime}$. An interleaving $a$-maximal cut says the cut made maximal progress for address $a$.
@@ 4.2 Side Accounts
Next, the FastSet validators will process this block as if it is issued from account ALICE_SIDE, which cause them to verify that $m$ satisfies verify({alice},1). Since $m$ is signed by Alice, it satisfies the specified quorum. FastSet validators will transfer 1 token from ALICE_SIDE to bob if the balance is sufficient.
@@ 4.3.2 Using a Template
The three mult-sig accounts ALICE_ABC, BOB_ABC, and CHARLIE_ABC created above have identical behaviors as the ones created in the previous sub-section.
@@ 4.3.4 A Dynamic Multi-Sig Contract
A more flexible contract for multisigs would allow an authority to establish and dynamically modify the signers and the quorum, while the signers would only sign. This can be done by first creating the following template:
@@
multisig_alice.do(transfer(gov,3); transfer(mama,1))alicesubscriptdelimited-⟨⟩monospace-$multisig_alice.do(transfer(gov,3); transfer(mama,1))|\rangle_{\texttt{alice}}$
@@
Some may argue that the above is not a problem, but a feature of FastSet. Indeed, it may very well be the case that david (and/or his proxy) was not malicious but was simply disconnected right before broadcasting the validator quorum. Implementations of FastSet may choose to also include the validator aggregated signature together with the message in $v.\textit{presettled}$ in Step 6 of the protocol, and then recheck the validity of verify in Step 7. We choose to not do it in setl, for efficiency reasons, but recommend developers and their auditors to think through such scenarios like above and handle them programmatically as desired. For example, the script can require instances to also provide the current signers and their quorum, and then check them against the contract’s:
@@ 4.4 Kudos
and then submits kudos_alice_bob.pay(3)bobsubscriptdelimited-⟨⟩monospace-$kudos_alice_bob.pay(3)|\rangle_{\texttt{bob}}$.
@@ 4.6 Settling Programming Languages and VMs
Many claims are expected to have a computational nature, such as “This Python program fibonacci on input 10 evaluates to 55”. In order for such a claim to be adequately verified, Python itself needs to be agreed upon, that is, settled as a verifiable artifact. The most rigorous such artifact would nevertheless be a formal semantics, paired with a series of expert security and design audits as well as formal properties proved about it using proof assistants and mechanical theorem provers (such as deterministic execution, type safety, etc.). While we truly hope that the future Web3 applications and users will demand such higher levels of rigor, the truth is that the current state of the art is far from it and it would be unreasonable to require such standards. Additionally, this would happen outside of FastSet anyway, and it ultimately boils down to the same basic principle of verifiers signing claims they are entitled to sign.
@@ 4.7 Settling Programs
where EVM is the unique ID/name of the claim that settles the specific EVM version used in this program. Both PYTHON and EVM can be settled as shown in Section 4.6. Same like PL or VM claims, program claims can and should also be set up to have verifiers, e.g., security (or ideally formal verification) auditors.
As we will shortly discuss in Section 4.8, program execution claims consist of a field holding the program claim, e.g., FIBONACCI or UNISWAP above, as well as the input to the program and the states before and after the execution. Clients can make multiple such computational claims and are responsible for properly composing them in order to avoid divergence (or double spending). However, such computational claims need to be verified. The client only proposes the computational claims, but cannot guarantee their integrity. In fact, the client may generate the claim using a cheap untrusted device, or even no device at all if the result can be inferred differently or guessed. Or can use powerful machines and parallel infrastructure that generates hundreds of computational claims per second, like some Layer 2 blockchains do.
@@ 4.8 Verifiable Computing
Despite its very limited direct programmability, FastSet makes it easy for clients to bring their own programming languages, virtual machines, and any computational models. All such languages do, from FastSet’s perspective, is to generate claims that need to be verified. It is client’s business to set their contracts up appropriately, to establish conventions for what languages are, who are the allowed verifiers, and what counts as a proof for them. We here show some possible ways to do it; sophisticated clients will likely find more creative solutions.
The following parametric template is parametric in a programming language LANG, a set of verifiers VERIFIERS, and a quorum of verifiers QUORUM.
@@
We are aware of the abuse of syntax in the above when we create the contract VERIFIED_EVM, where the parameter EVM_VERIFIERS should be used not as a value, but as an address. One must dereference EVM_VERIFIERS and create the contract using the actual value {V1,V2,...,V100000}. In this design document, we decide to leave with this abuse of syntax and treat it in a separate document where we nail down a formal specification of the setl language and its implementations.
@@
An application, such as a layer-2 blockchain L2chain, can now create an instance L2ver standing for “L2 verified”:
@@
Here, $T_{1}$ and $T_{2}$ are two (possibly different) sets of 1000 verifiers in EVM_VERIFIERS. The actual claims that FastSet validators will settle upon receiving the above messages, would be the following computational claims:
@@
Computational claims like above are meaningless for FastSet, but their settlement indicate that the verifiers approved them. Note that the verification of computational claims and calls into the L2ver.compute blocks take place in the verifiers, and thus outside FastSet.
@@
  • proof - a proof of the computation
Note that s and s’ are states of the program, and are not states of the FastSet validators. All the arguments except claims are self-explanatory. The claims argument corresponds to the claims that the application generates as the intended byproduct of the execution of program pgm, besides modifying its state. In fact, these claims should be considered part of the semantics of the program by the verifiers, and applications will likely provide specialized SDKs to their devs to generate these claims programmatically. For example, they can include payments (i.e., transfer claims) that the application determined to make while executing pgm, in particular to its verifiers who might not sign otherwise.
@@
The isSettled command checks if a claim has been settled. Every FastSet validator maintains a settled claim set. New claims, once they are validated and settled, will be added to this set. But no claims can be removed from the set. This way, the settled claim set forms an CRDT (see Section 2.2) w.r.t. claim settlement. In addition, isSettled is a side-effect free Boolean check on the monotonically increasing settled claim set, so once it is valid, it continues to stay valid. It thus guarantees that using isSettled to build a conditional claim does not break the weak independence property.
We glossed over many details here, such as how the programs and their inputs are represented, whether the states are given in clear or as indexes in an agreed upon DA layer, even over what constitutes a proof and how applications generate the claim arguments first place in order to present them to the verifiers. All these are FastSet-external conventions which are expected to be agreed upon in advance among the participating stakeholders (language committees/standards, applications, execution service providers, provers, verifiers, etc.).
@@ 4.9 Voting
Voting is an inherently parallel activity in which the order of votes is irrelevant, making FastSet ideal for it. An authority needs to set up and create a vote poll and then allow voters who are verified to vote. Finally, the authority should announce the results.
Let us consider the simplest voting scenario, where the authority, say gov, who is also responsible for verifying the voters, wants to accumulate MIN valid votes to pass some MOTION. Then gov can create a contract using the script:
@@
        not(instance.owner in voted);
@@
We’d like, however, to take the opportunity to further discuss this simple voting contract below, because it touches upon a few important aspects of FastSet that make it different from blockchains.
@@
Finally, it is insightful to note that although it may appear that a (malicious or not) voter can attempt to create two different instances and submit them concurrently to the validators in a hope that they both get settled, that is not possible thanks to the fact that each account has a nonce and at most one message is allowed in a validator’s pending for any given account (steps 3 and 4 in Figure 1). A voter attempting to create two concurrent (same nonce) instances would therefore risk to get their account stuck. There is nothing to prevent them to attempt to create two instances in sequence, each issued with a different nonce, but in that case they are processed in order and the second one will fail on all validators as soon as they processed the first one, as intended.
@@ 4.12 Non-Native Tokens and Digital Assets
circle_wallet.transfer_token(alice,1000)alicesubscriptdelimited-⟨⟩monospace-$circle_wallet.transfer_token(alice,1000)|\rangle_{\texttt{alice}}$
@@
All the non-native tokens and assets are stored in the contract, in its balance field. Whenever bob likes to get access to the 100 tokens that alice transfered to him, he can simply create an instance:
@@
and see that 100 tokens in his balance.
A successful token can result in many transfers made by many users in parallel, especially if used for payments. For example, VISA averages 20,000 transactions per second. With the advancement of AI and AI agents, it is expected that demand for higher rates for payments and micro-payments will significantly grow, perhaps to the order of millions of transfers per second. FastSet is ready for the challenge. It can settle a theoretically unlimited number of claims per second, including transfer claims like above. Indeed, there is nothing to prevent two transfers made by different accounts to proceed concurrently—we assume that location updates are atomic, which can be ensured using conventional synchronization mechanisms (e.g., software locks or hardware transactions).
@@
That implies v1 > 0 and balance[a1] >= v1, and respectively, v2 > 0 and balance[a2] >= v2. We have to prove that the balances of a1, a2, b1, and b2 are, respectively, the same no matter in which order the two transfer claims are processed. We analyze several cases, depending on whether b1 and b2 are equal or not, or if any of them or both are equal to any of a1 or a2. The most common case is that b1 and b2 are distinct and also distinct from a1 and a2. Then in both cases the balances are clearly the same: a1-v1, a2-v2, b1+v1, b2+v2. If b1==b2==b and are different from a1 and a2, then the balances are also the same: a1-v1, a2-v2, and b+v1+v2. If b1==a2 and b2 is distinct, then the balances are also the same: a1-v1, a2-v2+v1, and b2+v2. If b1==a2 and b2==a1, then the balances are the same, too: a1-v1+v2, a2-v2+v1. Finally, if b1==b2==a2 then the balances are: a1-v1, a2+v1. The remaining cases are similar. Consequently, the TOKEN contract above satisfies the weak independence requirement that guarantees its determinism no matter how the transfer claim blocks by different accounts are permuted. Because the (atomic) addition and subtraction operations on integers are commutative, an even stronger property holds: the individual claims in different blocks can also be further interleaved. This gives validators the freedom to maximize the parallelism, and thus the throughput, of token transfers.

### v3

@@ 3 Formalization and Correctness
However, the concurrency model underlying FastSet differs from classic (replica-based) distributed systems models like those in [21]. In distributed systems models, a number of communicating nodes perform a computation together. Accesses to shared objects impose certain constraints on the distributed computation, such as all the reads seeing the latest write value, but are also opportunities for nodes to share causality information. In FastSet, validators do not initiate any computation nor exchange any information with each other, neither directly nor indirectly through clients, and there is no causal dependence shared through updates (nor defined or calculated). Clients drive the protocol by broadcasting claims that may update the validator’s states. The only means of communication in FastSet, if it can even be called so, is that a client claim reaching a validator may be invalid for that validator, in which case the validator either discards it (in Step 3) or postpones it (in Step 6). An abstraction similar to our weak independence was introduced in [21] (Definition 2.6, Commutativity) and used to prove the SEC of their model, but it cannot be directly used in our formalization.
@@
Validators will never get stuck due to unfortunate interleaving orderings of claims. That is, if progress has been made by any other validators for any clients, then the validator in question can also make progress.
@@ 3.2 Weak Independence
Weak equivalence is reminiscent of the classic notion of “weak equality” in partial algebra (see, e.g., [4] for a survey): two claim sequences are weakly equivalent iff they result in the same state whenever they are both valid. What may be counter-intuitive at first sight is that an invalid claim sequence is weakly equivalent to any other claim sequence, whether valid or not—this meaning of “weak” is, however, well-established in the literature. This implies, in particular, that weak equivalence is not transitive: if $\gamma\Downarrow\gamma^{\prime}$ and $\gamma^{\prime}\Downarrow\gamma^{\prime\prime}$ then there is nothing to prevent the existence of a state $s$ such that ${\gamma}\!\downarrow_{s}$ and ${\gamma^{\prime\prime}}\!\downarrow_{s}$, but ${\gamma^{\prime}}\!\ndownarrow_{s}$ and $\llbracket{\gamma}\rrbracket{s}\neq\llbracket{\gamma^{\prime\prime}}\rrbracket{s}$. For example, $s$ can be a state where Alice has 11 tokens, $\gamma$ and $\gamma^{\prime\prime}$ payment claims by which Alice pays Bob 10 tokens and 11 tokens, respectively, and $\gamma^{\prime}$ some invalid claim (e.g., Alice pays Charlie 100 tokens).
@@
The reverse is not true. For example, if the two claims are the same mutex acquires, then each of them is valid independently, but they are not valid together regardless of the order. So $c\,c^{\prime}\Downarrow c^{\prime}\,c$ vacuously holds, but $c\parallel c^{\prime}$ does not hold.
@@ 3.3 Interleavings
Interleaving equivalence is reminiscent of actor/thread interleaving in concurrent programming, each address acting as an actor/thread executing a sequence of commands/claims. An interleaving cut $\gamma\sqsubseteq\gamma^{\prime}$ captures the intuition that $\gamma$ and $\gamma^{\prime}$ are two attempts to execute determined sequences of claims for each address, where $\gamma$ was cut short or did not advance as much as $\gamma^{\prime}$. An interleaving $a$-maximal cut says the cut made maximal progress for address $a$.
@@ 4 Examples and Applications
We remind the reader that the most common blockchain functionality, that of a payment system over a native token (e.g., Bitcoin), is a native support, as explained in Section 2.2. setl provides a primitive transfer(to,value) for this; e.g., if alice wants to transfer 3 native tokens to bob, then she submits and settles the claim transfer(bob,3)alicesubscriptdelimited-⟨⟩monospace-$transfer(bob,3)|\rangle_{\texttt{alice}}$.
@@ 4.2 Side Accounts
Next, the FastSet validators will process this block as if it is issued from account ALICE_SIDE, which causes them to verify that $m$ satisfies verify({alice},1). Since $m$ is signed by Alice, it satisfies the specified quorum. FastSet validators will transfer 1 token from ALICE_SIDE to bob if the balance is sufficient.
@@ 4.3.2 Using a Template
The three multi-sig accounts ALICE_ABC, BOB_ABC, and CHARLIE_ABC created above have identical behaviors as the ones created in the previous sub-section.
@@ 4.3.4 A Dynamic Multi-Sig Contract
A more flexible contract for multi-sigs would allow an authority to establish and dynamically modify the signers and the quorum, while the signers would only sign. This can be done by first creating the following template:
@@
multisig_alice.do(transfer(gov,3); transfer(mom,1))alicesubscriptdelimited-⟨⟩monospace-$multisig_alice.do(transfer(gov,3); transfer(mom,1))|\rangle_{\texttt{alice}}$
@@
Some may argue that the above is not a downside, but a feature of FastSet. Indeed, it may very well be the case that david (and/or his proxy) was not malicious but was simply disconnected right before broadcasting the validator quorum. Implementations of FastSet may choose to also include the validator aggregated signature together with the message in $v.\textit{presettled}$ in Step 6 of the protocol, and then recheck the validity of verify in Step 7. We choose to not do it in setl, for efficiency reasons, but recommend developers and their auditors to think through such scenarios like above and handle them programmatically as desired. For example, the script can require instances to also provide the current signers and their quorum, and then check them against the contract’s:
@@ 4.4 Kudos
and then submits kudos_alice_bob.pay(3)bobsubscriptdelimited-⟨⟩monospace-$kudos_alice_bob.pay(3)|\rangle_{\texttt{bob}}$. As a result, 3 native tokens are transferred from kudos_alice_bob to alice. Note that all accounts, and not just user-driven accounts, can hold native tokens. To ensure that the instance kudos_alice_bob has sufficient tokens, bob (or anyone else) should first transfer some tokens to it.
@@ 4.6 Settling Programming Languages and VMs
Many claims are expected to have a computational nature, such as “This Python program fibonacci on input 10 evaluates to 55”. In order for such a claim to be adequately verified, Python itself needs to be agreed upon, that is, settled as a verifiable artifact. The most rigorous such artifact would nevertheless be a formal semantics, paired with a series of expert security and design audits as well as formal properties proved about it using proof assistants and mechanical theorem provers (such as deterministic execution, type safety, etc.). While we truly hope that the future Web3 applications and users will demand such higher levels of rigor, the truth is that the current state-of-the-art is far from it and it would be unreasonable to require such standards just now. Additionally, this would happen outside of FastSet anyway, and it ultimately boils down to the same basic principle of verifiers signing claims they are entitled to sign.
@@ 4.7 Settling Programs
where EVM is the unique ID/name of the claim that settles the specific EVM version used in this program. Both PYTHON and EVM can be settled as shown in Section 4.6. Similar to PL/VM claims, program claims can and should also be set up to have verifiers, e.g., security (or ideally formal verification) auditors.
As we will shortly discuss in Section 4.8, program execution claims consist of a field holding the program claim, e.g., FIBONACCI or UNISWAP above, as well as the input to the program and the states before and after the execution. Clients can make multiple such computational claims and are responsible for properly composing them in order to avoid divergence (or double spending). However, such computational claims need to be verified. The client only proposes the computational claims, but cannot guarantee their integrity. In fact, the client may generate the claim using a cheap untrusted device, or even no device at all if the result can be inferred differently or guessed. Or can use powerful machines and parallel infrastructure that generates hundreds of computational claims per second, like some Layer 2 (L2) blockchains do.
@@ 4.8 Verifiable Computing
Despite its very limited direct programmability, FastSet makes it easy for clients to bring their own programming languages, virtual machines, and any computational models. All such languages do, from FastSet’s perspective, is to generate claims that need to be verified. It is the client’s responsibility to set their contracts up appropriately, to establish conventions for what languages are, who are the allowed verifiers, and what counts as a proof for them. We here show some possible ways to do it; sophisticated clients will likely find more creative solutions.
The following template is parametric in a programming language LANG, a set of verifiers VERIFIERS, and a quorum of verifiers QUORUM.
@@
We are aware of the abuse of syntax in the above when we create the contract VERIFIED_EVM, where the parameter EVM_VERIFIERS should be used not as a value, but as an address. One must dereference EVM_VERIFIERS and create the contract using the actual value {V1,V2,...,V100000}. In this design document, we decide to live with this abuse of syntax and treat it in a separate document where we nail down a formal specification of the setl language and its implementations.
@@
An application, such as a Layer 2 blockchain L2chain, can now create an instance L2ver standing for “L2 verified”:
@@
Here, $T_{1}$ and $T_{2}$ are two (possibly different) sets of 1000 verifiers in EVM_VERIFIERS. The actual claims that FastSet validators will settle upon receiving the above messages would be the following:
@@
Computational claims like above are meaningless for FastSet, but their settlement indicates that the verifiers approved them. Note that the verification of computational claims and calls into the L2ver.compute blocks take place in the verifiers, and thus outside FastSet.
@@
  • proof - a proof of the computation.
Note that s and s’ are states of the program, and not of the FastSet validators. All the arguments except claims are self-explanatory. The claims argument corresponds to the claims that the application generates as the intended byproduct of the execution of program pgm, besides modifying its state. In fact, these claims should be considered part of the semantics of the program by the verifiers, and applications will likely provide specialized SDKs to their devs to generate these claims programmatically. For example, they can include payments (i.e., transfer claims) that the application determined to make while executing pgm, in particular to its verifiers who might not sign otherwise.
@@
The isSettled command checks if a claim has been settled. Every FastSet validator maintains a settled claim set. New claims, once they are validated and settled, will be added to this set, but no claims can be removed from the set. In this way, the settled claim set forms an CRDT (see Section 2.2) w.r.t. claim settlement. In addition, isSettled is a side-effect free Boolean check on the monotonically increasing settled claim set, so once it is valid, it continues to stay valid. It thus guarantees that using isSettled to build a conditional claim does not break the weak independence property.
We glossed over many details here, such as how the programs and their inputs are represented, whether the states are given in clear or as indexes in an agreed upon data availability (DA) layer, even over what constitutes a proof and how applications generate the claim arguments first place in order to present them to the verifiers. All these are FastSet-external conventions which are expected to be agreed upon in advance among the participating stakeholders (language committees/standards, applications, execution service providers, provers, verifiers, etc.).
@@ 4.9 Voting
Voting is common on blockchains, yet it is an inherently parallel activity in which the order of votes is irrelevant, making FastSet ideal for it. An authority needs to set up and create a vote poll and then allow voters who are verified to vote. Finally, the authority should announce the results.
Let us consider a simple voting scenario, where the authority, say gov, who is also responsible for verifying the voters, wants to accumulate MIN valid votes to pass some MOTION. Then gov can create a contract using the setl script:
@@
        not(instance.owner in voted); // guard: must be true
@@
We would like to take the opportunity to further discuss this simple voting contract below, because it touches upon a few important aspects of FastSet that make it different from blockchains.
@@
Finally, it is insightful to note that although it may appear that a (malicious or not) voter can attempt to create two different instances and submit them concurrently to the validators in the hope that they both get settled, that is not possible thanks to the fact that each account has a nonce and at most one message is allowed in a validator’s pending for any given account (steps 3 and 4 in Figure 1). A voter attempting to create two concurrent (same nonce) instances would therefore risk to get their account stuck. A voter attempting to create two concurrent (same nonce) instances would therefore risk to get its account stuck. There is nothing to prevent them to attempt to create two instances in sequence, each issued with a different nonce, but in that case they are processed in order and the second one will fail on all validators as soon as they processed the first one, as intended. There is nothing to prevent voters from attempting to create two instances in sequence, each issued with a different nonce, but in that case they are processed in order and the second one will fail on all validators as soon as they processed the first one, as intended.
@@ 4.12 Non-Native Tokens and Digital Assets
circle_wallet.transfer_token(alice,1000)circlesubscriptdelimited-⟨⟩monospace-$circle_wallet.transfer_token(alice,1000)|\rangle_{\texttt{circle}}$
@@
All the non-native tokens and assets are stored in the contract, in its balance field. Whenever bob wants to get access to the 100 tokens that alice transferred to him, he can simply create an instance:
@@
and see the 100 tokens in his balance.
A successful token can result in many transfers made by many users in parallel, especially if used for payments. For example, VISA averages 20,000 transactions per second. With the advancement of AI and AI agents, it is expected that the demand for higher rates for payments and micro-payments will significantly grow, perhaps to the order of millions of transfers per second. FastSet is ready for the challenge. It can settle a theoretically unlimited number of claims per second, including transfer claims like above. Indeed, there is nothing to prevent two transfers made by different accounts to proceed concurrently—we assume that location updates are atomic, which can be ensured using conventional synchronization mechanisms (e.g., software locks or hardware transactions).
@@
That implies v1 > 0 and balance[a1] >= v1, and respectively, v2 > 0 and balance[a2] >= v2. We have to prove that the balances of a1, a2, b1, and b2 are, respectively, the same no matter in which order the two transfer claims are processed. We analyze several cases, depending on whether b1 and b2 are equal or not, or if any of them or both are equal to any of a1 or a2. The most common case is that b1 and b2 are distinct and also distinct from a1 and a2. Then in both cases the balances are clearly the same: a1-v1, a2-v2, b1+v1, b2+v2. If b1==b2==b and are different from a1 and a2, then the balances are also the same: a1-v1, a2-v2, and b+v1+v2. If b1==a2 and b2 is distinct, then the balances are also the same: a1-v1, a2-v2+v1, and b2+v2. If b1==a2 and b2==a1, then the balances are the same, too: a1-v1+v2, a2-v2+v1. Finally, if b1==b2==a2 then the balances are: a1-v1, a2+v1. The remaining cases are similar. Consequently, the TOKEN contract above satisfies the weak independence requirement that guarantees its determinism no matter how the transfer claim blocks by different accounts are permuted. Because the (atomic) addition and subtraction operations on integers are commutative, an even stronger property holds: the individual claims in different blocks can also be further interleaved. This gives validators the freedom to maximize the parallelism and thus the throughput of token transfers.